# Clusterpath Gaussian Graphical Modeling

### v2

@@ 5.1 S&P 100 Stocks
Since the data are temporally dependent, we first—as a preprocessing step—fit the popular heterogeneous autoregressive (HAR) model of Corsi, (2009) to the individual daily realized ranges to capture time dependencies, and then apply CGGM and TAGL to the standardized residuals to learn the conditional dependency structure among the stocks. See Appendix D.1 for details on data preprocessing and tuning parameter selection, including the selection of candidate values for $\phi$ based on the resulting weight distributions.
@@ 5.2 OECD Well-Being Indicators
The second data set we analyze is one on OECD well-being indicators, collected in 2018 by Cavicchia et al., (2022) (see Appendix D.2). The data contain $p=11$ variables related to well-being: education, jobs, income, safety, health, environment, civic engagement, accessibility to services, housing, community, and life satisfaction, on which two groups of countries with sample sizes $n_{1}=21$ and $n_{2}=14$ are given a score ranging from 0 to 10.
We highlight the capabilities of CGGM in estimating a cluster hierarchy on the standardized well-being data. Since our main goal is the retrieve the cluster hierarchy, we set $\lambda_{s}=0$ and report the whole clusterpath solution obtained by applying CGGM to the precision matrix with different values of the tuning parameter $\lambda_{c}$ for the two country groups. Due to the small group sample sizes, we do not use cross-validation to select the tuning parameters $\phi$ and $k$ but use fixed values based on existing conventions, namely $\phi=0.5$ and $k=3$ (e.g., Chi & Lange,, 2015; Wang et al.,, 2018). Since our primary interest is in the clusterpaths, the refitting step described in Section 2.3.4 is not required as it does not affect the obtained paths. Figure 8 visualizes the dendrograms as obtained from CGGM’s complete clusterpath solution ranging from $p=11$ clusters to one. If selection of a single solution along the clusterpath is requested, we advice practitioners to resort to domain expertise instead of data-driven methods when the sample size is that limited.
@@
The analysis of the well-being indicators across the two groups of countries reveals distinct clustering patterns that highlight differences in their socioeconomic conditions. In the first group (e.g., Australia, France, Switzerland), the aggregation of civic engagement, life satisfaction, community, environment, housing, accessibility to services, and jobs appears to reflect the social and environmental dimensions of well-being. Conversely, in the second group (e.g., Chile, Hungary, Mexico), safety and health are combined with the variables representing social and environmental dimensions, possibly suggesting a greater reliance on the community for these services. This contrasts with more developed countries where such services may be perceived as more self-evident. Additionally, in the second group, income, jobs, and education are grouped together, reflecting their importance to socioeconomic progress. Finally, note that we use CGGM with the precision matrix as object of interest; its block structure directly transfers to the covariance matrix thereby making our results comparable to those of Cavicchia et al., (2022) who estimate the clustering structure on the covariance matrix.
@@ 5.3 Humor Styles Questionnaire
In behavioral science surveys, multiple rating-scale items are typically used to measure each latent construct of interest. For further analysis, mean scores across the respective items are commonly computed as measurements of each latent construct, which implies the assumption that the covariance matrix of the questionnaire items follows a block structure. The humor styles questionnaire (HSQ) (Martin et al.,, 2003) measures four latent constructs corresponding to different styles of humor: affiliative, self-enhancing, aggressive, and self-defeating, with each construct being measured by eight items ($p=32$). If we can retrieve these four clusters of items in the covariance matrix, the assumption behind computing mean scores can be considered reasonable. Details on the HSQ data and tuning parameter selection in CGGM for clustering the covariance matrix are given in Appendix D.3.
CGGM returns four clusters of items that largely overlap with the grouping of Martin et al., (2003), as shown in Table 1. Looking at the wording of the questionnaire items, the clustering found by CGGM is intuitive. All items of the affiliative group, which CGGM places in Cluster 1, refer to humor in social settings. In addition, the two items of the self-enhancing group that complete Cluster 1—items 6 and 30—address being alone as opposed to being with people. On the other hand, the remaining six items of the self-enhancing group, which CGGM puts in Cluster 2, all refer to using humor for dealing with feeling depressed, upset, unhappy, sad, or having problems. Item 28 from the self-defeating group, which completes Cluster 2, also addresses having problems or feeling unhappy. In contrast, the other seven items of this group, which form Cluster 4 in the CGGM solution, refer to laughing or letting others laugh at oneself. Finally, all items from the aggressive group address teasing and offensive or inappropriate humor, and are placed in Cluster 3.
@@
Regarding sparsity, Martin et al., (2003, Table 2) report significant positive correlations between most of the four humor styles, at least for male participants. CGGM on the given data, on the other hand, finds nonzero covariance only between Clusters 1 and 2 as well as Clusters 2 and 4. That is, through regularization, the structure found by CGGM suggests latent variables that are more distinct than the factors found by Martin et al., (2003), which is desirable when characterizing different humor styles.
@@ 6 Conclusion
We introduce a novel method to estimate Gaussian graphical models (GGMs) subject to node-clustering in addition to edge-sparsity. Our method, which we call CGGM, uses a clusterpath penalty to produce a hierarchical clustering of variables (nodes in the graph) without relying on pre-existing notions of the cluster composition. Our software package CGGMR for the statistical computing environment R (R Core Team,, 2024) implements the proposed method and is available from https://github.com/djwtouw/CGGMR.
@@
Concerning future work, CGGM may be extended to partial correlation matrices instead of precision matrices. In contrast to the graphical lasso (GL), Carter et al., (2024) apply the sparsity penalty to the elements of the partial correlation matrix and call their method the partial correlation GL. As a result, the sparsity structure estimated by this method is invariant to scalar multiplication of the data. CGGM can also be modified to accommodate clustering based on the partial correlation matrix, but this comes at the cost of the convexity of the objective function as with the partial correlation GL.
@@
All computations were performed with R (R Core Team,, 2024). Replication files will be made publicly available upon acceptance and a link will be included here.
@@ A.1 Re-expressing the Objective Function $L(\mathbf{\Theta})$
$-\log|\bm{\Theta}|=-\log|\mathbf{P}^{1/2}\mathbf{R}^{*}\mathbf{P}^{1/2}|-\sum_{k=1}^{K}(p_{k}-1)\log a_{kk}.$
@@
$\sum_{j=1}^{p}\sum_{j^{\prime}=1}^{j-1}w_{jj^{\prime}}d_{jj^{\prime}}(\bm{\Theta})=\sum_{k=1}^{K}\sum_{\ell=1}^{k-1}\sum_{j\in\mathcal{C}_{k}}\sum_{j^{\prime}\in\mathcal{C}_{\ell}}w_{jj^{\prime}}d_{jj^{\prime}}(\mathbf{\Theta}),$
@@
where $\mathbf{Z}$ is the symmetric $p\times p$ weight matrix containing the sparsity weights $z_{jj^{\prime}}$.
@@
[…]
@@ A.2 Re-expressing the Objective Function $L(\mathbf{\Theta}_{k})$
Consequently, updating all parameters in $\mathbf{A}$ and $\mathbf{R}$ that are associated with cluster $k$ updates an entire block of variables in $\mathbf{\Theta}$. To separate the term $-\log|\mathbf{P}^{1/2}\mathbf{R}^{*}\mathbf{P}^{1/2}|$ in objective function (11), it helps to write $\mathbf{P}^{1/2}\mathbf{R}^{*}\mathbf{P}^{1/2}$ as the $2\times 2$ block matrix
@@
with $\mathbf{r}_{k}$ the $K-1$ vector containing all elements in the $k^{\text{th}}$ column of $\mathbf{R}$ apart from $r_{kk}$, and
@@
$-\log|\mathbf{P}^{1/2}\mathbf{R}^{*}\mathbf{P}^{1/2}|=-\log|\mathbf{P}_{0}^{1/2}\mathbf{R}_{0}^{*}\mathbf{P}_{0}^{1/2}|-\log(a_{kk}+p_{k}r_{kk}-p_{k}\mathbf{r}_{k}^{\top}(\mathbf{R}_{0}^{*})^{-1}\mathbf{r}_{k}).$ (12)
@@ B.1 Pseudocode for the Algorithm
Input Initial estimates for $\mathbf{B}^{(1)}$ and $\mathbf{R}^{(1)}$, sample covariance matrix $\mathbf{S}$, weight matrix       $\mathbf{W}$, tuning parameters $\lambda_{c}$ and $\lambda_{s}$, fusion threshold $\varepsilon_{f}$, convergence threshold $\varepsilon_{c}$, maxi-       mum number of iterations $t_{\text{max}}$
@@
1:$\mathbf{U}^{(1)}\leftarrow\mathbf{I}$
2:$L^{(1)}\leftarrow L(\mathbf{B}^{(1)},\mathbf{R}^{(1)})$
3:$L^{(0)}\leftarrow(1+2\varepsilon_{c})L^{(1)}$
4:$K\leftarrow p$
5:$t\leftarrow 1$
@@
21:$\widehat{\mathbf{B}}\leftarrow\mathbf{B}^{(t)}$
22:$\widehat{\mathbf{R}}\leftarrow\mathbf{R}^{(t)}$
23:$\hat{K}\leftarrow K$
24:$\widehat{\mathbf{A}}\leftarrow\text{diag}(\hat{a}_{11}\mathbf{I},\ldots,\hat{a}_{KK}\mathbf{I})$ where $\hat{a}_{kk}=\hat{b}_{kk}-\hat{r}_{kk}$
25:$\widehat{\mathbf{U}}\leftarrow\mathbf{U}^{(t)}$
26:$\widehat{\mathbf{\Theta}}\leftarrow\widehat{\mathbf{U}}\widehat{\mathbf{R}}\widehat{\mathbf{U}}^{\top}+\widehat{\mathbf{A}}$
@@
The algorithm requires initialization; the usage of high-quality initializations is important for the stability of the algorithm in terms of convergence. To initialize $\mathbf{B}$ and $\mathbf{R}$ for $\lambda_{c}=\lambda_{s}=0$, one can use the inverse of the sample covariance matrix if it is available; then $\mathbf{B}^{(1)}=\text{diag}(\mathbf{S}^{-1})$ and $\mathbf{R}^{(1)}=\mathbf{S}^{-1}$. In other cases, one can start from $(\mathbf{S}+\mathbf{I})^{-1}$ or a regularized estimator such as the graphical lasso (e.g., Friedman et al.,, 2008; Witten et al.,, 2011). Next, when computing the clusterpath, namely a sequence of solutions for $\mathbf{\Theta}$ for an increasing clustering parameter $\lambda_{c}$ (for fixed $\lambda_{s}$), we leverage warm starts to ensure high-quality initializations. Indeed, after solving the minimization problem for a particular $\lambda_{c}$, we use that solution as initialization when solving the optimization problem for the next (higher) value of $\lambda_{c}$ in the clusterpath. Loss function convergence of the algorithm was observed across all numerical experiments and empirical applications.
@@
After the required initializations, the algorithm first checks for eligible clusters fusions using the fusion threshold $\varepsilon_{f}$ (lines 12–15). The value of $\varepsilon_{f}$ can be a small user-defined value, such as $10^{-3}$, or based on the data itself. If the inverse of the sample covariance matrix is available, a data-driven choice is $\varepsilon_{f}=\tau\operatorname*{median}_{j,j^{\prime}}(d_{jj^{\prime}}(\mathbf{S}^{-1}))$, with $\tau=10^{-3}$. If a fusion is performed, the optimization parameters are modified and the total number of clusters $K$ is reduced by one. In case no cluster fusion occurs, the algorithm proceeds to the second step where the parameters pertaining to cluster $k$, denoted by the vector $[b_{kk},\mathbf{r}_{k},r_{kk}]$, are jointly updated (lines 17–20). Part of the update is a line search by means of a golden section search for the optimal step size $s^{*}$, which ensures the positive definiteness of $\mathbf{\Theta}$.
@@
As mentioned in Appendix A, $\mathbf{\Theta}$ is positive definite if $\mathbf{R}^{*}$ is positive definite and $a_{kk}=b_{kk}+r_{kk}>0$. Assume that $\mathbf{R}^{*}$ is positive definite for the current values of $[b_{kk},\mathbf{r}_{k},r_{kk}]$. Then, Sylvester’s Criterion states that the upper left $1\times 1$ corner of $\mathbf{R}^{*}$ has a positive determinant, so does the $2\times 2$ upper left corner, and so on until $\mathbf{R}^{*}$ itself has a positive determinant. Without loss of generality, let the clusters be arranged so that $k$ is the last cluster. Updating $[b_{kk},\mathbf{r}_{k},r_{kk}]$ does not alter the $1\times 1$ through $(K-1)\times(K-1)$ upper left corners. Consequently, a positive determinant of $\mathbf{R}^{*}$ after the update of the parameters related to cluster $k$ is sufficient to keep $\mathbf{R}^{*}$ positive definite.
@@
After the loss function has converged, the output of the algorithm consists of the estimated number of clusters $\hat{K}$ and the estimates $\widehat{\mathbf{B}}$, $\widehat{\mathbf{R}}$, and $\widehat{\mathbf{U}}$, which can be used to construct $\widehat{\mathbf{\Theta}}$.
The computational complexity of the algorithm is primarily determined by the computation of the descent direction $\bm{\delta}_{k}$, which requires solving a system of equations with a complexity of $\mathcal{O}(K^{3})$. This step dominates the overall update cost, including the line search. To determine $s^{*}$, we use the golden section search, which iteratively shrinks the interval $[0,s_{\max}]$ to a pre-specified tolerance of $5\cdot 10^{-3}$. During this search, the objective for cluster $k$ in (13) is repeatedly evaluated, incurring a per-evaluation complexity of $\mathcal{O}(K^{2})$. Consequently, a complete pass over all clusters results in a total complexity of $\mathcal{O}(K^{4})$. Since the number of clusters $K$ is bounded by the number of variables $p$, the overall complexity is capped at $\mathcal{O}(p^{4})$.

### v3

@@ 5.1 S&P 100 Stocks
Since the data are temporally dependent, we first—as a preprocessing step—fit the popular heterogeneous autoregressive (HAR) model of \@BBOPcite\@BAP\@BBNCorsi (2009)\@BBCP to the individual daily realized ranges to capture time dependencies, and then apply CGGM and TAGL to the standardized residuals to learn the conditional dependency structure among the stocks. See Appendix D.1 for details on data preprocessing and tuning parameter selection, including the selection of candidate values for $\phi$ based on the resulting weight distributions.
@@ 5.2 OECD Well-Being Indicators
The second data set we analyze is one on OECD well-being indicators, collected in 2018 by \@BBOPcite\@BAP\@BBNCavicchia et al. (2022)\@BBCP (see Appendix D.2). The data contain $p=11$ variables related to well-being: education, jobs, income, safety, health, environment, civic engagement, accessibility to services, housing, community, and life satisfaction, on which two groups of countries with sample sizes $n_{1}=21$ and $n_{2}=14$ are given a score ranging from 0 to 10.
We highlight the capabilities of CGGM in estimating a cluster hierarchy on the standardized well-being data. Since our main goal is the retrieve the cluster hierarchy, we set $\lambda_{s}=0$ and report the whole clusterpath solution obtained by applying CGGM to the precision matrix with different values of the tuning parameter $\lambda_{c}$ for the two country groups. Due to the small group sample sizes, we do not use cross-validation to select the tuning parameters $\phi$ and $k$ but use fixed values based on existing conventions, namely $\phi=0.5$ and $k=3$ (e.g., \@BBOPcitep\@BAP\@BBNChi and Lange, 2015; Wang et al., 2018\@BBCP). Since our primary interest is in the clusterpaths, the refitting step described in Section 2.3.4 is not required as it does not affect the obtained paths. Figure 8 visualizes the dendrograms as obtained from CGGM’s complete clusterpath solution ranging from $p=11$ clusters to one. If selection of a single solution along the clusterpath is requested, we advice practitioners to resort to domain expertise instead of data-driven methods when the sample size is that limited.
@@
The analysis of the well-being indicators across the two groups of countries reveals distinct clustering patterns that highlight differences in their socioeconomic conditions. In the first group (e.g., Australia, France, Switzerland), the aggregation of civic engagement, life satisfaction, community, environment, housing, accessibility to services, and jobs appears to reflect the social and environmental dimensions of well-being. Conversely, in the second group (e.g., Chile, Hungary, Mexico), safety and health are combined with the variables representing social and environmental dimensions, possibly suggesting a greater reliance on the community for these services. This contrasts with more developed countries where such services may be perceived as more self-evident. Additionally, in the second group, income, jobs, and education are grouped together, reflecting their importance to socioeconomic progress. Finally, note that we use CGGM with the precision matrix as object of interest; its block structure directly transfers to the covariance matrix thereby making our results comparable to those of \@BBOPcite\@BAP\@BBNCavicchia et al. (2022)\@BBCP who estimate the clustering structure on the covariance matrix.
@@ 5.3 Humor Styles Questionnaire
In behavioral science surveys, multiple rating-scale items are typically used to measure each latent construct of interest. For further analysis, mean scores across the respective items are commonly computed as measurements of each latent construct, which implies the assumption that the covariance matrix of the questionnaire items follows a block structure. The humor styles questionnaire (HSQ) \@BBOPcitep\@BAP\@BBN(Martin et al., 2003)\@BBCP measures four latent constructs corresponding to different styles of humor: affiliative, self-enhancing, aggressive, and self-defeating, with each construct being measured by eight items ($p=32$). If we can retrieve these four clusters of items in the covariance matrix, the assumption behind computing mean scores can be considered reasonable. Details on the HSQ data and tuning parameter selection in CGGM for clustering the covariance matrix are given in Appendix D.3.
CGGM returns four clusters of items that largely overlap with the grouping of \@BBOPcitet\@BAP\@BBNMartin et al. (2003)\@BBCP, as shown in Table 1. Looking at the wording of the questionnaire items, the clustering found by CGGM is intuitive. All items of the affiliative group, which CGGM places in Cluster 1, refer to humor in social settings. In addition, the two items of the self-enhancing group that complete Cluster 1—items 6 and 30—address being alone as opposed to being with people. On the other hand, the remaining six items of the self-enhancing group, which CGGM puts in Cluster 2, all refer to using humor for dealing with feeling depressed, upset, unhappy, sad, or having problems. Item 28 from the self-defeating group, which completes Cluster 2, also addresses having problems or feeling unhappy. In contrast, the other seven items of this group, which form Cluster 4 in the CGGM solution, refer to laughing or letting others laugh at oneself. Finally, all items from the aggressive group address teasing and offensive or inappropriate humor, and are placed in Cluster 3.
@@
Regarding sparsity, \@BBOPcitet\@BAP\@BBNMartin et al. (2003, Table 2)\@BBCP report significant positive correlations between most of the four humor styles, at least for male participants. CGGM on the given data, on the other hand, finds nonzero covariance only between Clusters 1 and 2 as well as Clusters 2 and 4. That is, through regularization, the structure found by CGGM suggests latent variables that are more distinct than the factors found by \@BBOPcitet\@BAP\@BBNMartin et al. (2003)\@BBCP, which is desirable when characterizing different humor styles.
@@ 6 Conclusion
[…]
@@
Concerning future work, CGGM may be extended to partial correlation matrices instead of precision matrices. In contrast to the graphical lasso (GL), \@BBOPcite\@BAP\@BBNCarter et al. (2024)\@BBCP apply the sparsity penalty to the elements of the partial correlation matrix and call their method the partial correlation GL. As a result, the sparsity structure estimated by this method is invariant to scalar multiplication of the data. CGGM can also be modified to accommodate clustering based on the partial correlation matrix, but this comes at the cost of the convexity of the objective function as with the partial correlation GL.
@@
[…]
@@ A.1 Re-expressing the Objective Function $L(\mathbf{\Theta})$
$\log|\bm{\Theta}|=-\log|\mathbf{P}^{1/2}\mathbf{R}^{*}\mathbf{P}^{1/2}|-\sum_{k=1}^{K}(p_{k}-1)\log a_{kk}.$
@@
$\sum_{j=2}^{p}\sum_{j^{\prime}=1}^{j-1}w_{jj^{\prime}}d_{jj^{\prime}}(\bm{\Theta})=\sum_{k=1}^{K}\sum_{\ell=1}^{k-1}\sum_{j\in\mathcal{C}_{k}}\sum_{j^{\prime}\in\mathcal{C}_{\ell}}w_{jj^{\prime}}d_{jj^{\prime}}(\mathbf{\Theta}),$
@@
where $\mathbf{Z}$ is the symmetric $p\times p$ weight matrix containing the sparsity weights $z_{jj^{\prime}}$. We take the sparsity weights $z_{jj^{\prime}}$ to be the absolute value of the corresponding entry in $\mathbf{S}^{-1}$ or $(\mathbf{S}+\mathbf{I})^{-1}$ when the inverse of the sample covariance matrix is not available.
@@
$={}$ $-\log|\mathbf{P}^{1/2}\mathbf{R}^{*}\mathbf{P}^{1/2}|-\sum_{k=1}^{K}(p_{k}-1)\log a_{kk}+\operatorname*{tr}\mathbf{SURU}^{\top}$
@@ A.2 Re-expressing the Objective Function $L(\mathbf{\Theta}_{k})$
Consequently, updating all parameters in $\mathbf{A}$ and $\mathbf{R}$ that are associated with cluster $k$ updates an entire block of variables in $\mathbf{\Theta}$. To separate the term $\log|\mathbf{P}^{1/2}\mathbf{R}^{*}\mathbf{P}^{1/2}|$ in objective function (11), it helps to write $\mathbf{P}^{1/2}\mathbf{R}^{*}\mathbf{P}^{1/2}$ as the $2\times 2$ block matrix
@@
with $\mathbf{r}_{k}$ the K1K-1 vector containing all elements in the $k^{\text{th}}$ column of $\mathbf{R}$ apart from $r_{kk}$, and
@@
$\log|\mathbf{P}^{1/2}\mathbf{R}^{*}\mathbf{P}^{1/2}|=-\log|\mathbf{P}_{0}^{1/2}\mathbf{R}_{0}^{*}\mathbf{P}_{0}^{1/2}|-\log(a_{kk}+p_{k}r_{kk}-p_{k}\mathbf{r}_{k}^{\top}(\mathbf{R}_{0}^{*})^{-1}\mathbf{r}_{k}).$ (12)
@@ B.1 Pseudocode for the Algorithm
Input Initial estimates for $\mathbf{B}^{(1)}$ and $\mathbf{R}^{(1)}$, sample covariance matrix $\mathbf{S}$, weight matrix    $\mathbf{W}$, tuning parameters $\lambda_{c}$ and $\lambda_{s}$, thresholds $\varepsilon_{f}$ and $\varepsilon_{s}$, convergence threshold $\varepsilon_{c}$, maxi-    mum number of iterations $t_{\text{max}}$
@@
[…]
@@
[…]
@@
The algorithm requires initialization; the usage of high-quality initializations is important for the stability of the algorithm in terms of convergence. To initialize $\mathbf{B}$ and $\mathbf{R}$ for $\lambda_{c}=\lambda_{s}=0$, one can use the inverse of the sample covariance matrix if it is available; then $\mathbf{B}^{(1)}=\text{diag}(\mathbf{S}^{-1})$ and $\mathbf{R}^{(1)}=\mathbf{S}^{-1}$. In other cases, one can start from $(\mathbf{S}+\mathbf{I})^{-1}$ or a regularized estimator such as the graphical lasso (e.g., \@BBOPcitep\@BAP\@BBNFriedman et al., 2008; Witten et al., 2011\@BBCP). Next, when computing the clusterpath, namely a sequence of solutions for $\mathbf{\Theta}$ for an increasing clustering parameter $\lambda_{c}$ (for fixed $\lambda_{s}$), we leverage warm starts to ensure high-quality initializations. Indeed, after solving the minimization problem for a particular $\lambda_{c}$, we use that solution as initialization when solving the optimization problem for the next (higher) value of $\lambda_{c}$ in the clusterpath. Loss function convergence of the algorithm was observed across all numerical experiments and empirical applications.
@@
After the required initializations, the algorithm first checks for eligible clusters fusions using the fusion threshold $\varepsilon_{f}$ (lines 12–15). The value of $\varepsilon_{f}$ can be a small user-defined value, such as $10^{-3}$, or based on the data itself. If the inverse of the sample covariance matrix is available, a data-driven choice is $\varepsilon_{f}=\tau\operatorname*{median}_{j,j^{\prime}}(d_{jj^{\prime}}(\mathbf{S}^{-1}))$, with $\tau=10^{-3}$. If a fusion is performed, the optimization parameters are modified and the total number of clusters $K$ is reduced by one. In case no cluster fusion occurs, the algorithm proceeds to the second step where the parameters pertaining to cluster $k$, denoted by the vector $b_{kk},\mathbf{r}_{k},r_{kk}]$, are jointly updated (lines 17–20). Part of the update is a line search by means of a golden section search for the optimal step size $s^{*}$, which ensures the positive definiteness of $\mathbf{\Theta}$.
@@
As mentioned in Appendix A, $\mathbf{\Theta}$ is positive definite if $\mathbf{R}^{*}$ is positive definite and $a_{kk}=b_{kk}+r_{kk}>0$. Assume that $\mathbf{R}^{*}$ is positive definite for the current values of $b_{kk},\mathbf{r}_{k},r_{kk}]$. Then, Sylvester’s Criterion states that the upper left $1\times 1$ corner of $\mathbf{R}^{*}$ has a positive determinant, so does the $2\times 2$ upper left corner, and so on until $\mathbf{R}^{*}$ itself has a positive determinant. Without loss of generality, let the clusters be arranged so that $k$ is the last cluster. Updating $b_{kk},\mathbf{r}_{k},r_{kk}]$ does not alter the $1\times 1$ through $(K-1)\times(K-1)$ upper left corners. Consequently, a positive determinant of $\mathbf{R}^{*}$ after the update of the parameters related to cluster $k$ is sufficient to keep $\mathbf{R}^{*}$ positive definite.
@@
After the loss function has converged, a thresholding operation is applied to $\mathbf{R}$ to set sufficiently small elements to zero and induce sparsity. Finally, the output of the algorithm consists of the estimated number of clusters $\hat{K}$ and the estimates $\widehat{\mathbf{B}}$, $\widehat{\mathbf{R}}$, and $\widehat{\mathbf{U}}$, which can be used to construct $\widehat{\mathbf{\Theta}}$.
The computational complexity of the algorithm is primarily determined by the computation of the descent direction $\bm{\delta}_{k}$, which requires solving a system of equations with a complexity of $\mathcal{O}(K^{3})$. This step dominates the overall update cost, including the line search. To determine $s^{*}$, we use the golden section search, which iteratively shrinks the interval $0,s_{\max}]$ to a pre-specified tolerance of $5\cdot 10^{-3}$. During this search, the objective for cluster $k$ in (13) is repeatedly evaluated, incurring a per-evaluation complexity of $\mathcal{O}(K^{2})$. Consequently, a complete pass over all clusters results in a total complexity of $\mathcal{O}(K^{4})$. Since the number of clusters $K$ is bounded by the number of variables $p$, the overall complexity is capped at $\mathcal{O}(p^{4})$.